OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 6 − x = 5x + 30 4x − 8 = 3x + 2 The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describe solutions to equations shown.

OpenStudy (anonymous):

6-x=5x+30 divide -1 on both sides 6=-5x-30 Add 30 on both sides 36=-5x Divide by -5 on both sides\[x=-7.2\] 4x-8=3x+2 Add 8 on both sides 4x=3x+10 Subtract 3x from both sides \[x=10\] THE SOLUTION TO EQUATION #1 IS SMALLER THAN THE SOLUTION TO EQUATION #2